8-(3b+5)=-5+2(b+1)
-8n+5>-2n-12
x+3/5 +x/3=7
3p-p/4-5=p/6-2p+6
4x+3(x-2)=18-x
These are some problems I've been working on which have me puzzled.

Answer by rfer(16322)

8-(3b+5)=-5+2(b+1)
8-3b-5=-5+2b+2
-5b=-6
b=-6/5
b=-1 1/5
------------
-8n+5 > -2n-12
-6n > -17
n > -17/-6
n > 2 5/6
----------------
x+3/5+x/3=7
2x/3=7-3/5
2x/3=35/5-3/5
2x/3=32/5
2x=32/5*3
2x=96/5
x=96/5*1/2
x=96/10
x=9.6